For the given sample data and null hypothesis, compute the value of the test statistic, z

A claim is made that the proportion of children who play sports is less than 0.5, and the sample statistics include 1933 subjects with 30% saying that they play a sport.

Question options:

A)

35.90

B)

0.50

C)

-17.59

D)

-35.90

E)

17.59
Answer #1

Answer Given that claim pis less than 0.5 Hypothesis can be written as Ho P=05 Hi-Pros P=05 and q=1-1=1-05:05 9-1933 9 - 301